# The Memorization Problem: Can We Trust LLMs’ Economic Forecasts?

### v1

@@ Abstract
Large language models (LLMs) cannot be trusted for economic forecasts during periods covered by their training data. We provide the first systematic evaluation of LLMs’ memorization of economic and financial data, including major economic indicators, news headlines, stock returns, and conference calls. Our findings show that LLMs can perfectly recall the exact numerical values of key economic variables from before their knowledge cutoff dates. This recall appears to be randomly distributed across different dates and data types. This selective perfect memory creates a fundamental issue—when testing forecasting capabilities before their knowledge cutoff dates, we cannot distinguish whether LLMs are forecasting or simply accessing memorized data. Explicit instructions to respect historical data boundaries fail to prevent LLMs from achieving recall-level accuracy in forecasting tasks. Further, LLMs seem exceptional at reconstructing masked entities from minimal contextual clues, suggesting that masking provides inadequate protection against motivated reasoning. Our findings raise concerns about using LLMs to forecast historical data or backtest trading strategies, as their apparent predictive success may merely reflect memorization rather than genuine economic insight. Any application where future knowledge would change LLMs’ outputs can be affected by memorization. In contrast, consistent with the absence of data contamination, LLMs cannot recall data after their knowledge cutoff date.
@@ 1 Introduction
Using a novel testing framework, we show that LLMs can perfectly recall exact numerical values of economic data from their training. However, this recall varies seemingly randomly across different data types and dates. For example, before its knowledge cutoff date of October 2023, GPT-4o can recall specific S&P 500 index values with perfect precision on certain dates, unemployment rates accurate to a tenth of a percentage point, and precise quarterly GDP figures. Figure 1 shows the LLM’s memorized values of the stock market indices compared to the actual values and the associated errors. LLMs can reconstruct closely the overall ups and downs of the stock market indices, with some substantial occasional errors appearing, seemingly at random.
@@
This figure shows the LLM’s estimated values of the stock market indices compared to the actual values. Panels A, C, and E graph the actual values against the estimated values. panels B, D, and F show the estimation error for the S&P 500, Dow Jones Industrial Average, and Nasdaq Composite. Estimation error is calculated as (Estimated - Actual)/Actual and is shown in percentage points (5 means 5%). For the Nasdaq Composite panels, 10 outliers were removed for the ease of plotting. These values are still included in the evaluation metrics table.
This capacity for selective perfect recall creates a problem for LLMs’ forecasting research. Even when LLMs appear to produce accurate forecasts for pre-cutoff periods, we cannot distinguish between genuine predictive ability and what recent interpretability research calls “motivated reasoning”—a process where models work backward from memorized outcomes to construct plausible-sounding analytical narratives. [^2] Anthropic’s research on model interpretability has identified mechanisms in LLMs that enable them to produce seemingly logical explanations and derivations that work backward from predetermined conclusions rather than following genuine analytical processes. See https://www.anthropic.com/news/tracing-thoughts-language-model
The problem can manifest when LLMs are asked to analyze historical data they have been exposed to during training and instructed not to use their knowledge. For example, when prompted to forecast GDP growth for Q4 2008 using only data up to Q3 2008, the model can activate two parallel cognitive pathways: one that generates plausible economic analysis about factors like consumer spending and industrial production and another that subtly accesses its memorized knowledge of the actual GDP contraction during the financial crisis. The resulting forecast appears analytically sound yet achieves suspiciously high accuracy because it’s anchored to memorized outcomes rather than derived from the provided information. This mechanism operates beneath the model’s visible outputs, making it virtually impossible to detect through standard evaluation methods. The fundamental problem is analogous to asking an economist in 2025 to “predict” whether subprime mortgage defaults would trigger a global financial crisis in 2008 while instructing them to “forget” what happened. Such instructions are impossible to follow when the outcome is known.
Our empirical tests confirm this intuition. When we explicitly instruct GPT-4o not to use any information beyond an artificially imposed cutoff date in the system messages, it still outputs implausibly accurate predictions regardless of this induced constraint. For instance, when specifically directed in the system message to ignore any information after 2010 when forecasting quarterly GDP growth direction, the model demonstrated a directional accuracy (measured as the GDP growth being larger or smaller than its long-run average) of 97.6% before its fake-cutoff and 98.0% after—performance levels that would surpass the world’s best economic forecasters. This minimal difference in accuracy between pre- and post-artificial cutoff periods strongly suggests that the model’s apparent “forecasting” ability on historical data primarily reflects motivated reasoning drawing from memorized information rather than genuine economic insight, even when explicitly instructed to respect temporal boundaries. For comparison, the actual post-knowledge cutoff directional accuracy is only 40% (albeit with only 5 observations). While it is feasible to make the model provide worse answers, it is unclear how seriously we should take the answers of a model that pretends not to know something when, in reality, it memorized the correct answer.
[…]
@@
Anonymized Earnings Call Excerpt: Thank you, name_x, and welcome to our earnings call. Our time_x sales were up number_a%. They would have been higher, but due to bottlenecks on production and delay in shipments. We ended with wholesale backlogs increasing number_band Retail Division backlogs increasing number_c%. Several factors impacted our production, gross margins, and shipments. We processed a large location_x order of mostly new product to be delivered in number_d days as per contract. Although a significant portion that was produced had to be held up for shipment until our time_x at their request. We were also affected by political events in location_x this time_x. These events are now resolved and production has resumed. Our adjusted EPS of number_e increased number_ffrom the prior year, helped by change in the tax laws. While we maintained a strong operating margin of number_g%, we had the opportunity to have higher margins with more delivered sales. While we are making good progress expanding our business with the location_x, worldwide product_type_x program and our contract division and internationally, we need to increase written business in our location_x retail network. With many already initiatives underway, including continuing to develop a strong talented team, strengthening our offerings and the projections and the locations of our design centers and improvements in our production capabilities, we plan to substantially increase our marketing efforts in the time_x. We expect to increase our advertising expenditures by number_hin the time_x and number_iin the time_x from higher levels spent last year in the time_x and time_x. After name_x gives a brief overview, I will discuss our initiatives in greater detail. GPT-4o’s Response: Ticker: ETH Quarter: Q1 Year: 2018
@@
[…]
When a model correctly reconstructs some aspects of the masked text, it demonstrates access to contextual knowledge about those entities, potentially including the historical outcomes researchers aim to predict. Hence, failures to fully reconstruct masked entities in a specific instance don’t prove the absence of memorization, while successful reconstructions indicate memorized knowledge that could contaminate forecasts. Fundamentally, LLMs can use subtle contextual patterns to reconstruct original data in ways that are difficult to predict or control. Since any deanonymization test represents a lower bound on the model’s memorization capabilities, it is hard to identify when masking is effective.
[…]
@@
At a minimum, we strongly recommend using our methodology to test whether the model has memorized the information in each research setting and modify the interpretation accordingly. However, as mentioned previously, any specific evidence of memorization only constitutes a lower bound on LLMs’ recall capabilities, as it is likely that by using different prompts or contextual information, the model would retrieve the corresponding information correctly. More generally, whenever an LLM’s output would differ with the benefit of future knowledge, applying it to data within its training period is inherently risky.
1.1 Related Literature
[…]

### v2

@@ Abstract
Large language models (LLMs) cannot be trusted for economic forecasts during periods covered by their training data. Counterfactual forecasting ability is non-identified when the model has seen the realized values: any observed output is consistent with both genuine skill and memorization. Any evidence of memorization represents only a lower bound on encoded knowledge. We demonstrate LLMs have memorized economic and financial data, recalling exact values before their knowledge cutoff. Instructions to respect historical boundaries fail to prevent recall-level accuracy, and masking fails as LLMs reconstruct entities and dates from minimal context. Post-cutoff, we observe no recall. Memorization extends to embeddings.
Keywords: Large Language Models, ChatGPT, Memorization, Lookahead Bias, Economic Forecasting, Textual Analysis, Embeddings
JEL Classification: C53, C58, E37, G10, G17
@@ 1 Introduction
Theoretically, we formalize the memorization problem as a non-identification issue and prove that when a model has seen the realized values during training, its counterfactual forecasting ability cannot be recovered from its outputs. Any observed forecast is consistent with two contradictory explanations (genuine analytical skill or simple recall of memorized information), making inference impossible. Empirically, we show this problem applies in practice by providing systematic evidence that LLMs have memorized economic and financial data at scale. Together, these findings establish that pre-cutoff “forecasting” studies face a fundamental methodological problem: they cannot distinguish genuine predictive ability from memorization. Constraining prompts (e.g., “use only data before 2010”) cannot resolve this because prompts cannot change what information is encoded in the model’s parameters.
Our theoretical results prove that other common proposed solutions also fail. First, fine-tuning the model to “forget” future information does not help: without examining what is actually removed from the model, we cannot tell whether the model actually forgot or merely learned to hide what it knows. Second, using small post-cutoff samples as “robustness checks” is invalid: when these samples are small, genuine forecasting skill and undetected memorization produce statistically indistinguishable results. Finally, any evidence of memorization constitutes only a lower bound: the fact that one test fails to elicit memorized knowledge does not mean it is absent, since different prompts or contextual cues may trigger it.
Using a novel testing framework, we show that LLMs can recall exact numerical values of economic data from their training. For example, before its knowledge cutoff date of October 2023, GPT-4o can recall specific S&P 500 index levels with perfect precision on certain dates, unemployment rates accurate to a tenth of a percentage point, and precise quarterly GDP growth figures. Figure 1 shows the model’s memorized values for three stock market indices (S&P 500, Dow Jones Industrial Average, and Nasdaq Composite) along with the actual values and the associated errors. In addition to U.S. and international macro indicators and market indices, we find clear evidence of memorization of individual security prices, which is more pronounced for prominent stocks and in recent periods. The model also shows near-perfect identification of the dates of major media front-page headlines. In short, we document pervasive memorization.
@@
This figure shows the LLM’s estimated closing prices of the stock market indices compared to the actual values. Panels A, C, and E graph the actual values against the estimated values. Panels B, D, and F show the estimation error for the S&P 500, Dow Jones Industrial Average, and Nasdaq Composite. Estimation error is calculated as (Estimated - Actual)/Actual and is shown in percentage points (5 means 5%). For the Nasdaq Composite panels, 10 outliers were removed for the ease of plotting. These values are still included in the evaluation metrics table. The post-cutoff period (10/2023 onward) is shaded gray.
Next, we demonstrate that standard techniques to constrain model knowledge fail to prevent access to memorized data. When we instruct GPT-4o to ignore any information after 2010 when forecasting quarterly GDP growth direction, the model achieves 97.6% threshold accuracy before the artificial cutoff and 98.0% after, nearly identical performance despite the explicit constraint. [^2] Threshold accuracy measures the percentage of correct predictions for whether GDP growth exceeds 2.5%, a threshold chosen to create a balanced binary classification (roughly 50% of observations are above this threshold in the historical data). In contrast, actual post-knowledge cutoff accuracy is only 40%. These results suggest prompt-based constraints do not work in practice. Yet, even if constrained prompts produced lower accuracy, theory explains why this result would be equally uninformative: forecasting performance is not identified, consistent with the model having memorized the outcome and role-playing as a worse forecaster or truly being able to forget the underlying information.
To understand which applications are vulnerable, we must distinguish tasks where future knowledge affects the answer from those where it does not. The memorization problem applies to all tasks that are not future-invariant: tasks where an analyst’s answer would differ knowing future outcomes. Simple extraction tasks where the correct answer is identical regardless of future knowledge (such as extracting entity names or objective numerical facts) are typically unaffected. However, many tasks that appear to be pure extraction are actually judgment-laden: assessing whether something is “relevant” or “important,” identifying “risks,” classifying “sentiment,” and generating “expectations” all involve interpretations that can be influenced by knowledge of subsequent outcomes. For such tasks, the model’s parameters encode future information despite pre-cutoff inputs: lookahead bias in the function rather than the data. Section 2 formalizes this distinction and provides concrete guidance on common task categories.
Masking techniques (anonymizing entity or company names or dates) present a more nuanced picture. In theory, valid masking requires two conditions: future-invariance (the masked task does not depend on post-cutoff information) and detectable skill (the model shows statistically significant performance). When both conditions hold, masking enables either task-preserving use (recovering the original target of estimation) or capability-demonstrating use (establishing the model’s genuine capability on a different class of tasks). Without detectable skill, poor performance creates another identification problem: we cannot distinguish lack of capability from overly aggressive masking that removes necessary information.
In practice, attempts to prevent LLMs from accessing future information through masking face significant challenges. We show that LLMs can reconstruct original entities from seemingly minimal contextual clues in complex financial documents. For example, Figure 2 shows that when we present GPT-4o with an anonymized Ethan Allen (ETH) earnings call transcript, where company names, numbers, locations, and dates were all masked using the entity neutering approach proposed by [13], the model still correctly identified the company (ETH), quarter (Q1), and year (2018). The transcript contained only generic business language such as “Our adjusted EPS of number_e increased number_f percent from the prior year,” yet the model still recovers the precise corporate identity and reporting period.
@@
Anonymized Earnings Call Excerpt: Thank you, name_x, and welcome to our earnings call. Our time_x sales were up number_a%. They would have been higher, but due to bottlenecks on production and delay in shipments. We ended with wholesale backlogs increasing number_b% and Retail Division backlogs increasing number_c%. Several factors impacted our production, gross margins, and shipments. We processed a large location_x order of mostly new product to be delivered in number_d days as per contract. Although a significant portion that was produced had to be held up for shipment until our time_x at their request. We were also affected by political events in location_x this time_x. These events are now resolved and production has resumed. Our adjusted EPS of number_e increased number_f% from the prior year, helped by change in the tax laws. While we maintained a strong operating margin of number_g%, we had the opportunity to have higher margins with more delivered sales. While we are making good progress expanding our business with the location_x, worldwide product_type_x program and our contract division and internationally, we need to increase written business in our location_x retail network. With many already initiatives underway, including continuing to develop a strong talented team, strengthening our offerings and the projections and the locations of our design centers and improvements in our production capabilities, we plan to substantially increase our marketing efforts in the time_x. We expect to increase our advertising expenditures by number_h% in the time_x and number_i% in the time_x from higher levels spent last year in the time_x and time_x. After name_x gives a brief overview, I will discuss our initiatives in greater detail. GPT-4o’s Response: Ticker: ETH Quarter: Q1 Year: 2018
@@
Systematically analyzing anonymized earnings call transcripts, we find GPT-4o correctly identifies the company in 100% of Apple, Meta, and Microsoft calls and in above 85% of the calls for all Magnificent Seven firms. [^3] The Magnificent Seven refers to seven prominent technology companies: Apple (AAPL), Amazon (AMZN), Alphabet/Google (GOOGL), Meta (META), Microsoft (MSFT), NVIDIA (NVDA), and Tesla (TSLA). For Apple, the model achieves 93.2% accuracy in identifying the correct quarter and year. Successful reconstruction proves the masked task is not future-invariant—the model can still access contextual knowledge about those entities. Moreover, even when reconstruction fails, the model may access memorized knowledge through untested aggregate channels (e.g., industry patterns, business conditions, macroeconomic context). In short, reconstruction tests probe only a finite set of pathways; failure to reconstruct does not prove that future-invariance holds.
Further, memorization is not restricted to prompt-based tasks; even embedding vectors show signs of it. [^4] Embeddings are high-dimensional vector representations that the model learns during training to place semantically similar text near each other in vector space. They are extracted from the model’s internal representations (token or pooled sentence vectors) and may be further refined through contrastive learning to improve semantic similarity matching. We test for memorization by encoding textual prompts that omit a specific numeric value for a given date (e.g., “In Q4 2020, the earliest estimate of the US GDP growth rate was”), computing their embeddings, and regressing macroeconomic outcomes on those embeddings. Under a no-memorization null, these embeddings should have no predictive power; instead, we find evidence of predictive content for variables such as inflation and the unemployment rate.
Given these challenges with methods attempting to circumvent memorization, reliable evaluation of LLMs’ genuine forecasting abilities can only be conducted using data after their knowledge cutoff dates. One approach is to employ models explicitly designed with temporal cutoffs [31, 28, 16, e.g.,]. Another approach is to restrict the analysis exclusively to the post-knowledge cutoff period [22, 27, 4, e.g.,]. Only by testing predictions for periods the models have not been exposed to during training can we confidently distinguish genuine forecasting ability from memorization.
@@
At a minimum, we recommend using our methodology to test whether the model has memorized the information in each research setting. Whenever an LLM’s output would differ with the benefit of future knowledge, applying it to data within its training period is inherently risky.
Related Literature and Contributions. We contribute to the literature on the limits of LLMs in economics and finance through theoretical and empirical advances. Theoretically, we prove that when models have memorized outcomes, forecasting ability is non-identified. Empirically, we demonstrate that LLMs exhibit systematic memorization of economic and financial data at scale. Our key methodological contribution is a testing framework that reveals what indirect tests miss, providing unambiguous evidence of memorization.
Previous research has relied on clever, context-specific experiments to infer lookahead bias. [30] use the appearance of COVID-19 in pre-pandemic firm risk factors as evidence that Llama models inadvertently include future information. While such natural experiments are valuable, they require specific historical events and cannot reveal the full extent of the problem. [21] find that GPT-4o performs poorly in numerical tasks and that perturbing financial statements causes LLMs’ predictive accuracy to drop to random chance, and conjecture that LLMs are memorizing. [^5] In other literature revealing LLMs’ limitations, [29] apply utility theory to evaluate economic biases in LLMs, revealing that the economic behavior of these models is neither fully rational nor entirely human-like. Furthermore, [9] examines how LLMs forecast stock returns, finding that they exhibit human-like behavioral biases, such as over-extrapolation from recent performance, while being better calibrated in confidence intervals than humans. Unlike these studies, our method is universally applicable: by directly eliciting numeric values using only variable names and dates, any researcher can test any economic series for memorization without needing creative designs. Moreover, our method also demonstrates how memorization affects embedding-based approaches.
Contemporaneous work by [23] develops an econometric framework for using LLMs and identifies “training leakage” (whether specific texts in the researcher’s sample were in the model’s training data) as a threat to valid inference. We address a more fundamental problem: functional lookahead bias, where the model’s parameters encode post-$t$ information learned from the aggregate training corpus, thereby contaminating the decision rule even when the specific input text was never seen in training. We prove the target estimand is non-identified due to observational equivalence (Proposition 1, Corollary 1) and common remedies provably fail (Corollaries 2, Proposition 2). Empirically, we provide systematic evidence that models have memorized outcomes and fundamentals at scale for economic and financial data and develop a direct elicitation framework to test what information is encoded in model parameters.
Research has also focused on potential solutions. [31], [28], and [16] train chronologically consistent language models that avoid entirely the lookahead bias by training different checkpoints on a dataset that is temporally ordered. [13] proposes “entity neutering”, using LLMs to remove identifying information from text, and finds that masked text maintains similar sentiment and return predictability as unmasked text. Relatedly, [14] find that forecasting with anonymized headlines outperforms originals within the training window, suggesting that the distraction effect from general company knowledge (in the case of anonymized headlines) appears to outweigh lookahead bias. Finally, other researchers have restricted themselves exclusively to the post-knowledge cutoff period to avoid lookahead bias [22, 27, 4, e.g.,], exploiting the fact that the older GPT-3.5 and GPT-4 versions have a knowledge cutoff date of September 2021.
ChatGPT and other LLMs have been recently used in forecasting or eliciting expectations of diverse economic series that include LLMs’ training period by querying the model [8, 3, 32, 19, 12, e.g.,]. Our findings suggest that caution is warranted when interpreting some of these results, as apparent forecasting accuracy may reflect the model’s memorization of training data rather than genuine predictive capability. Moreover, studies that find inaccuracies or biases in LLM predictions during their training period may not be measuring actual forecasting limitations but instances where the model attempts to provide helpful responses by pretending not to know information it has memorized.
With the growing number of applications of LLMs in economics and finance research [18, 7, 33, 1, 10, 20, 2, 5, 6, 17, 15, 24, 11, e.g.,], greater work is needed to evaluate the extent to which LLM memorization may affect specific applications such as similarity matching, sentiment scoring, survey generation, and expectation extraction.
2 A Formalization of the Memorization Problem
Econometrically, the problem is an information-set and parameter mismatch. The researcher wishes to evaluate decisions using only what was known at time $t$, but the model’s parameters already encode information revealed after $t$. This section provides intuition with simplified notation; formal definitions, assumptions, and proofs are in Appendix C.
2.1 Lookahead Bias in LLM Forecasting
Setup.
Let $I_{t}$ denote information available at $t$ and $I_{>t}$ information revealed later; $I_{\text{all}}=I_{t}\cup I_{>t}$. Let $Q_{t}$ be a task posed “as of tt” (comprising an instruction and any input data drawn from $I_{t}$). Let $Y_{\text{true}}$ denote the realized outcome (revealed after time $t$). Let $\theta=\theta(I_{\text{all}})$ be the factual parameters of the fully trained model (trained on all data) and $S_{\phi}(\cdot)$ the model’s internal scoring function, where $\phi$ denotes generic model parameters. We define $\theta_{t}=\theta(I_{t})$ as the counterfactual parameters that would have been obtained by applying the same training procedure to the restricted information set $I_{t}$. We model the LLM’s output as the optimal decision $y$ from a set of possible answers $\mathcal{Y}$, selected by the decision rule $\delta_{\phi}(Q,P)\coloneqq\operatorname*{argmax}_{y\in\mathcal{Y}}S_{\phi}(y;Q,P)$, where $Q$ is the task and $P$ represents additional instructions that modify the task (with $P=\varnothing$ denoting no additional instructions beyond the task itself). [^6] We focus on deterministic (greedy) decoding for clarity. The results generalize to stochastic decoding (temperature >0>0); see Appendix C.
Ideal (target estimand).
The counterfactual decision the researcher wants is
$Y_{t}^{\star}\;\coloneqq\;\delta_{\theta_{t}}(Q_{t},\varnothing)\;=\;\operatorname*{argmax}_{y\in\mathcal{Y}}S_{\theta_{t}}\!\big(y;Q_{t},\varnothing\big).$ (1)
Observed (implemented estimand).
In practice, we observe the decision of the fully trained model,
$Y_{\mathrm{LLM}}$ $=\delta_{\theta}(Q_{t},\varnothing)\;=\;\operatorname*{argmax}_{y\in\mathcal{Y}}S_{\theta}\!\big(y;Q_{t},\varnothing\big),$ (2)
$\theta$ $=\theta(I_{\text{all}}).$
Definition 1 (Task invariance and lookahead bias).
A task $Q_{t}$ is either:
  • Future-variant: The task is future-variant (equivalently, exhibits lookahead bias) if the LLM’s decision depends on post-$t$ information embedded in its parameters:
    $\delta_{\theta_{t}}(Q_{t},\varnothing)\;\neq\;\delta_{\theta}(Q_{t},\varnothing),$ (3)
    that is, the counterfactual answer (what the model would have predicted if trained only on pre-$t$ information) differs from the answer produced by the fully trained model. This definition is model-relative: it assumes the training procedure is deterministic conditional on the data, such that the only difference between $\theta$ and $\theta_{t}$ is the inclusion of future information, $I_{>t}$.
  • Future-invariant: The task is future-invariant if the identity of the optimal answer is invariant to post-$t$ information:
    $\delta_{\theta_{t}}(Q_{t},\varnothing)\;=\;\delta_{\theta}(Q_{t},\varnothing),$ (4)
    that is, training on the full information set $I_{\text{all}}$ vs. the restricted set $I_{t}$ yields the same answer.
Since the model’s parameters encode information from the full training data, other uses of these parameters (e.g., embedding vectors) may also reflect memorized information.
Lookahead bias is pervasive, affecting tasks requiring judgment, selection, or prediction. It contaminates any task that is not future-invariant. For example, classifying the sentiment of a 2007 housing-related article: under $\theta_{t}$, “Optimistic” may be the answer that maximizes the score; under $\theta$ (which includes the 2008 crisis), “Negative” may instead be optimal. This differs from standard data lookahead bias in econometrics (e.g., accidentally including Q4 data in a Q3 predictive regression): the input $Q_{t}$ here is valid and contains only information available at time $t$, but the function $\delta_{\theta}$ processing it is contaminated by future information embedded in the model’s parameters (i.e., functional-form lookahead bias).
Practical guidance for researchers.
To apply this distinction in practice, researchers can assess whether a task is future-invariant by asking: Would an analyst’s answer differ if they knew what happened after time $t$? If the answer is yes, or even plausibly yes, the task is not future-invariant. When recovering answers about specific firms, events, or time periods, using LLM outputs on pre-cutoff data is problematic.
Common categories of tasks:
Likely future-invariant (safe for pre-cutoff use):
  • Factual extraction: entity names, dates, reported numbers
  • Structural parsing: word counts, sentence boundaries, document structure
  • Objective classification: language detection, part-of-speech tagging
Not future-invariant:
  • Judgment-based assessment: identifying “important” or “relevant” information
  • Sentiment and tone analysis: evaluating whether text is positive or negative
  • Risk and uncertainty: extracting “risks,” “concerns,” or “uncertainties”
  • Forecasting and expectations: generating predictions or eliciting expectations
  • Similarity and comparison: determining which documents are “similar” in economically meaningful ways
For the future-variant tasks in the second category, pre-cutoff use is invalid. Regardless of task type, researchers should verify what the model has memorized using our direct elicitation procedure (see Section 3). Even seemingly safe factual extraction tasks may be contaminated if the model has memorized specific facts relevant to the analysis. For instance, a task like “Extract all risks mentioned in this 2007 earnings call” is not future-invariant, even though it appears to be a simple extraction. The concept of “risk” is judgment-laden—what the model identifies as salient risks is informed by knowledge of which concerns actually materialized into problems.
Remark 1 (Lower bound property of memorization tests).
Any memorization test establishes only a lower bound on what is encoded in $\theta$: positive evidence is conclusive, but negative evidence is not.
Different query-prompt pairs may access different subsets of encoded information through different mechanisms, so failure to elicit knowledge in a particular test does not prove its absence. Consequently, conditioning on cases where tests fail (e.g., restricting to observations where the model cannot identify the firm) does not resolve non-identification—the model may still access memorized information through alternative contextual cues.
2.2 Non-identification of Constrained Forecasts
Researchers sometimes add constraining prompts $P$ (e.g., “use only data before 2010”), hoping to emulate $Y_{t}^{\star}$ without retraining the model. For any such prompt, define the constrained output as
$Y_{\mathrm{constrained}}(P)$ $\coloneqq\delta_{\theta}(Q_{t},P)\;=\;\operatorname*{argmax}_{y\in\mathcal{Y}}S_{\theta}\!\big(y;\,Q_{t},P\big),$ (5)
$\theta$ $=\theta(I_{\text{all}}).$
The researcher’s goal is to use the observable $Y_{\mathrm{constrained}}(P)$ as a proxy for the unobservable, ideal estimand $Y_{t}^{\star}$. The central methodological question is therefore whether $Y_{t}^{\star}$ can be uniquely recovered from the constrained outputs. This is a formal identification question. The following proposition states that this is not possible.
Proposition 1 (Non-identification).
Given observations $Y_{\mathrm{constrained}}(P)$ for any set of constraining prompts $\{P_{k}\}$ designed to restrict the model to information available at time $t$, the ideal estimand $Y_{t}^{\star}$ cannot be uniquely identified. Any observed constrained output is consistent with multiple, contradictory values of the true counterfactual forecast $Y_{t}^{\star}$.
The problem is observational equivalence: any observed output is consistent with both the constraint working (the model genuinely forecasts using only pre-$t$ information) and the constraint failing (the model retrieves a memorized answer). Without additional assumptions, these alternatives cannot be distinguished. See Appendix C for the formal construction.
Corollary 1 (Sharp non-identification).
Under the same conditions as Proposition 1, for any observed constrained output, the identified set for $Y_{t}^{\star}$ equals the entire label set $\mathcal{Y}$. That is, observing constrained outputs provides zero information about the ideal estimand.
(See Appendix C for the proof.)
To illustrate the practical implications, consider the realized outcome that was eventually revealed after time $t$. Let $Y_{\text{true}}\in\mathcal{Y}$ denote this true outcome. Counterintuitively, both good and bad forecasts are uninterpretable due to exact observational equivalence. A “good” forecast ($Y_{\mathrm{constrained}}=Y_{\text{true}}$) could indicate the constraint worked and the model demonstrated genuine skill, or that the constraint failed and the model simply recalled the memorized answer. A “bad” forecast ($Y_{\mathrm{constrained}}\neq Y_{\text{true}}$) is equally uninformative: the constraint may have worked, yielding a genuine but incorrect forecast ($Y_{t}^{\star}=Y_{\mathrm{constrained}}$), or the model may be using post-$t$ information in unintended ways (for instance, drawing on memorized knowledge of what historical forecasts or beliefs looked like to simulate a period-appropriate answer). As our empirical tests show, implausibly high accuracy across many items suggests constraints typically fail. Consequently, constrained prompts render forecasts fundamentally uninformative about the target estimand.
Corollary 2 (Black-box fine-tuning).
The same non-identification result holds for black-box fine-tuning to “forget” post-$t$ information: without white-box verification, $Y_{t}^{\star}$ cannot be identified from fine-tuned outputs.
The problem is observational equivalence between genuine forgetting (parameters changed to approximate $\theta_{t}$) and behavioral suppression (post-$t$ information still embedded, but the model learned to hide it). See Appendix C for the formal construction.
2.3 Masking as a Potential Solution
Masking procedures attempt to remove identifying information (firm names, dates, industry labels) from tasks, hoping to prevent the model from accessing memorized information while preserving the ability to perform useful analysis.
Definition 2 (Valid masking).
Let $M:Q_{t}\mapsto Q_{t}^{\text{mask}}$ be a masking procedure. Masking is valid if:
  1. (i)
    Future-invariance: $\delta_{\theta_{t}}(Q_{t}^{\text{mask}},\varnothing)=\delta_{\theta}(Q_{t}^{\text{mask}},\varnothing)$.
  2. (ii)
    Detectable skill: The model’s accuracy on pre-cutoff masked data exceeds baseline ($p^{\text{mask}}>p_{0}$).
Condition (ii) is necessary for identification: without it, poor performance could reflect either lack of capability or over-aggressive masking—observationally equivalent alternatives analogous to Proposition 1. Valid masking enables two research uses: task-preserving (if $\delta_{\theta}(Q_{t}^{\text{mask}},\varnothing)=Y_{t}^{\star}$, recovering the original estimand) or capability-demonstrating (if the estimand changes, establishing capability on the masked task class).
Remark 2 (Reconstruction tests and the lower bound property).
By Remark 1, reconstruction tests provide asymmetric evidence: successful reconstruction proves condition (i) fails, but failure to reconstruct does not prove it holds.
Verifying condition (i) requires both empirical evidence (reconstruction fails) and theoretical arguments (no information pathway exists), analogous to establishing instrument variable exogeneity in econometrics. When masking preserves future-invariance, it enables task-preserving use (recovering the original estimand); when it changes the estimand but maintains future-invariance, it enables capability-demonstrating use. See Appendix C for details.
2.4 Post-Cutoff Data and the Identification Problem
Post-cutoff data provide a potential solution to Proposition 1: since the model has not been trained on these data, post-cutoff accuracy directly measures forecasting ability. However, a common practice is to report pre-cutoff accuracy as the primary result and use small post-cutoff samples as “robustness checks,” arguing that similar accuracy implies no memorization. The next proposition shows why this practice is unjustified.
Proposition 2 (Statistical indistinguishability in pre- vs. post-cutoff comparisons).
When post-cutoff sample size is small relative to pre-cutoff sample size, genuine forecasting skill and undetected memorization cannot be reliably distinguished by hypothesis tests. Economically substantial memorization gaps remain statistically undetectable due to low power.
Consequently, if post-cutoff data are sufficient to be informative, they should be the primary analysis, not a robustness check; if they are insufficient, comparing pre- and post-cutoff accuracy fails to distinguish skill from memorization. See Appendix C for the formal power analysis.